Discuss the market outcome of a tariff. How does it affect the market participants? Which participant gain and which participant lose? Explain all gains and loss using . ( NO PLAGIARISM PLEASE)
Answer #1

Impact of tariff (assuming small country) is analyzed below.

In the graph, AB & CD are domestic demand & supply curves of the concerned good.

In absence of trade, domestic equilibrium is at point E with price P1 & quantity Q1.

With free trade with zero tariff, world price is P*, against which, domestic consumption is Q2 and domestic production is Q3. So,

Imports = Q2 - Q3.

Consumer surplus (CS) = Area between demand curve and world price = Area of triangle AFP*

Producer surplus (PS) = Area between supply curve and world price = Area of triangle CGP*

An import tariff will increase the domestic price of the good from P* to Pt (with Pt - P* = unit tariff), against which domestic consumption is Q4 and domestic production is Q5.

Imports = Q4 - Q5.

So, the tariff decreases imports.

After tariff,

New CS = Area of triangle AHPt

New PS = Area of triangle CJPt.

Decrease in CS = Area PtHFP*, so tariff hurts consumers.

Increase in PS = Area PtJGP*, so tariff benefits domestic producers.

Tariff revenue = Area of rectangle PtHLP*, so tariff benefits the government.

Net Social loss = Area of triangle FHL + Area of triangle GJK

63 91 9 9, 82 8